sinx +siny =-21/65 and cosx +cosy =-27/65 then cos(x+y/2

Grade:11

1 Answers

Arun
5 years ago
Dear Rahul
 
(sinx+siny)= – 21/65
cosx+cosy= – 27/65
Dividing this we get
tan(x + y/2)=21/27
cos (x+y/2)=27/sqrt (1170)
 
Regards
Arun (askIITians forum expert)